For a normal distribution with a mean of 13 and a standard deviation of 6, the value –2 has a z value of
Answer by stanbon(75887)

For a normal distribution with a mean of 13 and a standard deviation of 6, the value –2 has a z value of
------------------
z(-2) = (-2-13)/6 = -15/6 = -2.5
----
Cheers,
Stan H.
=======